OpenStudy (anonymous):

A bacteria population experiences continuous growth. If there were 2,000 bacteria present at the start of the experiment, how many did it take for there to be 6,000 bacteria if the growth rate was 7% per hour? I will grant you a fan and a medal.

OpenStudy (misty1212):

hI!!

OpenStudy (misty1212):

start with the equation for growth and 7% per hour, do you know it?

OpenStudy (anonymous):

Hi!

OpenStudy (anonymous):

2,000 * 0.07?

OpenStudy (misty1212):

not quite

OpenStudy (misty1212):

to increase a number by \(7\%\) multiply it by \[100\%+7\%=107\%=1.07\]

OpenStudy (misty1212):

to do it say \(t\) times, multiply it by \(1.07^t\)

OpenStudy (anonymous):

OH! So were timesing it by 1.07, okay.

OpenStudy (misty1212):

right and your job is to solve \[2000\times (1.07)^t=6000\]any clues ?

OpenStudy (misty1212):

"no clue" is ok i can show you, just asking

OpenStudy (anonymous):

Yep, totally clueless.

OpenStudy (misty1212):

ok first step is obvious probably , divide by 2000

OpenStudy (misty1212):

\[1.07^t=3\]

OpenStudy (misty1212):

so far so good?

OpenStudy (misty1212):

did i lose you there?

OpenStudy (anonymous):

2.80?

OpenStudy (anonymous):

Is that what t is?

OpenStudy (misty1212):

i doubt it triples that fast

OpenStudy (misty1212):

you need either a fancy calculator or logs so solve this does your class do logs and exponents?

OpenStudy (anonymous):

Yes.

OpenStudy (misty1212):

are you familiar with the "change of base" formula"?

OpenStudy (anonymous):

No...

OpenStudy (misty1212):

\[b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\] the log of the total divided by the log of the base

OpenStudy (misty1212):

so \[1.07^t=3\iff t=\frac{\ln(3)}{\ln(1.07)}\] and a calculator

OpenStudy (misty1212):

what class has conic sections and logs? wowee

OpenStudy (anonymous):

Pre-Calc.

OpenStudy (misty1212):

really mixing it up did you get an answer ?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

16.24